OpenStudy (anonymous):

Find the time in hours and minutes between 22:30 and 4:20 the next day. (I have never been good at this...Help me plz!!! :D)

OpenStudy (thecurious):

22:30 - 23:30 = 60 minutes 23:30- 00:30 = 60 mins 00:30 - 1:30 = 60 mins 1:30 - 2:30 = 60 mins 2:30 - 3:30 = 60 mins 3:30 - 4:20 = 50 mins that means 5 hours 50 mins!

OpenStudy (anonymous):

Thank you!!!! Math has always been my worst subject...I have straight A's-with a B in math...ugh...